# Advanced Detection of Source Code Clones via an Ensemble of Unsupervised Similarity Measures

### v1

@@ Abstract
Knowledge Engineering, Similarity Measures, Code Similarity
@@ 1 Introduction
Automatically assessing the degree of similarity between source code fragments presents several challenges ranging from syntactic variability or the use of complex algorithmic structures to the employment of diverse abstraction levels across programming languages [18]. While established techniques and tools exist to address this problem, significant limitations must be addressed. For example, the best-performing approaches need vast training data, their insights are practically intelligible to the human operator, and their associated carbon footprint during the training phase does not bring positive societal outcomes [30].
To overcome these issues, this research proposes a novel ensemble approach for improved code similarity assessment. The method combines the outputs of multiple similarity measures to produce better similarity scores. This approach is based on the principle that the strengths of diverse similarity measures can complement each other and mitigate individual weaknesses, leading to improved systemic performance.
Extensive empirical evaluations on several datasets have demonstrated the effectiveness of our ensemble learning method. Our method outperformed traditional single-model approaches in all cases (both in scenarios with small and large datasets). Furthermore, our method achieved comparable performance to state-of-the-art methods when solving small datasets without reliance on deep learning techniques, making it a solution to consider under circumstances such as the absence of significant amounts of training data, the need for interpretability and the wish for a reduced carbon footprint when training.
Therefore, our contribution is not only to guide the aggregation of appropriate unsupervised similarity measures for code similarity but also to pave the way for applications such as source clone detection [35] or code plagiarism assessment [31] in realistic scenarios where there is not too much data for training. In this way, the major contributions of this study are:
  • We build upon previous work [28] to curate a comprehensive collection of unsupervised similarity measures to compare the likeness of different code fragments. This collection facilitates a systematic exploration of the most effective measures for assessing source code similarity at low computational cost.
@@
  • Empirical evaluation of the proposed ensembles focuses on key metrics such as precision, recall, and f-measure. Our findings suggest that ensembles outperform individual similarity measures and have a strong potential to rival even the most sophisticated state-of-the-art techniques in the absence of significant training data.
@@ 2 State-of-the-art
The existence of many programming languages and individual coding styles poses a challenge for determining the degree of likeness between source code fragments [1]. Despite that, unsupervised similarity measures demonstrate reasonable adaptability when working with code written in multiple languages or according to various styles, so it can be accepted that they are good indicators of code similarity [28]. This feature is important in many scenarios, such as maintenance, whereby increased complexity is commonplace [15].
The research community widely acknowledges that the emerging strategies based on language models offer superior performance in this context [19]. However, our hypothesis is that unsupervised methods for code similarity evaluation offer several advantages for addressing key aspects that should not be overlooked:
  • Code understanding: Capturing the meaning of code is a complex task, as functionally equivalent code fragments may exhibit superficial differences or, on the contrary, share surface-level similarities while having entirely different purposes. Some unsupervised techniques specifically focus on the semantic intent rather than syntactic details.
  • Independence from labeled data: Since unsupervised techniques eliminate the need for labeled training data, they reduce the overhead of establishing a ground truth. While labeled examples remain valuable for validating the effectiveness of these methods, these examples are not a prerequisite for deployment.
  • Noise filtering: Codebases often contain comments and other non-functional elements embedded within the core logic. Specific unsupervised methods can distinguish between meaningful code patterns and just incidental material.
  • Stylistic diversity: Unsupervised methods can accommodate the diversity of programming languages and coding styles. This mitigates the need for a complex strategy that attempts to account for all possible variations.
We acknowledge the superior capabilities of language models when confronted with this challenge. However, we also recognize and aim to demonstrate that alternative models, which excel in specific areas, could be employed under certain plausible conditions across a wide range of everyday scenarios. For instance, these models may be more suitable in situations where only limited datasets are available or where the working logic of the proposed solution needs to be comprehended.
@@ 2.1 Unsupervised Code Similarity
There are various techniques to assess the similarity between source code fragments, each with its unique perspective based on certain features of the code fragments being compared. The following is a list that we have already proposed previously [28], consisting of twenty-one distinct unsupervised approaches. Table LABEL:tab:long presents all these similarity measures.
Table 1: Comparison of Different Unsupervised Similarity Approaches
Approach Description Ref.
Abstract Syntax Trees (ASTs) Similarity Examines the hierarchical structures of code to identify structural similarities between different ASTs. [22]
Bag-of-Words Similarity Counts the occurrence of words in texts, comparing these frequencies while ignoring word order and structural context. [7]
Code Embeddings Similarity Uses vector representations of code for comparison, evaluating similarity based on these embeddings. This technique is used without recalibration (zero-shot) here. [2]
Comments Similarity Focuses on comparing code comments to facilitate documentation and understanding, using traditional text similarity methods. [29]
Fuzzy Matching Similarity Adapts string comparison techniques for minor syntactical variations to code similarity assessment. [37]
Function Calls Similarity Compares source code fragments based on their functions names. [42]
Graph-based Similarity Relies on the relationships in graphs, representing data structures or code dependencies, to determine similarity. [44]
Jaccard Similarity Uses set intersection and union to evaluate similarity in text analysis and information retrieval contexts. [13]
Levenshtein Similarity Measures the minimum number of edits needed to transform one string into another, reflecting similarity. [25]
Longest Common Subsequence (LCS) Similarity Identifies the longest sequence of items that two series share, indicating similarity. [5]
Metrics Similarity Involves computing various code-related metrics and comparing these values to estimate similarity. [32]
N-grams Similarity Compares texts based on the occurrence of contiguous sequences of n items, evaluating shared n-grams. [8]
Output Analysis Similarity Assesses the similarity of program outputs by applying traditional text similarity metrics. [29]
Perceptual Hashing Similarity Adapts techniques from image similarity, using hash codes from visual representations of code to measure similarity. [33]
Program Dependence Graph Similarity Assesses the similarity between codes by analyzing their program dependence graphs, which represent the dependencies between their elements. [24]
Rolling Hash Similarity Employs a dynamically updating hash to compare substrings in large texts, adapted for code comparison. [14]
RKR-GST Similarity Targets code similarity detection by identifying maximal sequences of contiguous matching tokens in texts. [41]
Semdiff Similarity Detects semantic differences between program versions, evaluating the impact of code changes on semantics. [16]
Semantic Clone Similarity Assesses code fragment similarity based on the semantic meanings of program element names. [11]
TF-IDF Similarity Common in text analysis, compares texts based on the relative importance of terms, using weighted term frequencies. [20]
Winnow Similarity Identifies similar texts by hashing and comparing text fingerprints. [36]
According to existing research [28], it is unrealistic to think that these techniques alone can deliver performance equivalent to the state-of-the-art. However, we aim to show how their aggregation can do so under certain conditions.
@@ 2.2 Supervised Code Similarity
Supervised code similarity is a well-studied field and has yielded the best results to date [38, 39, 40, 45, 43]. Although among the existing approaches, CodeBERT [10] and its variant GraphCodeBERT [12] stand out. CodeBERT is a variant of BERT [9], and empirical results show that it is the leading technology to understand source code expressed in different general-purpose programming languages. It excels at tasks like generating documentation, creating code, and comparing similar pieces of code. The idea behind CodeBERT is to combine techniques from language processing with an understanding of code and training on a wide range of programming languages and their associated comments. Results show CodeBERT is very good at detecting similarities between source code fragments, as demonstrated by its performance in the resolution of benchmark datasets for clone detection [10]. However, an additional fine-tuning phase on the pre-trained model over the actual data is usually required to obtain optimal results.
@@ 2.3 Contribution over the state-of-the-art
This research introduces a novel ensemble learning approach for code similarity assessment. The rationale behind combining multiple unsupervised similarity measures is to leverage the strengths of individual approaches while mitigating their weaknesses. This strategy is helpful in scenarios where deep learning is not optimal due to the absence of significant training data, when prioritizing model interpretability without significantly compromising performance, or even when seeking to minimize the environmental impact.
@@ 3 Ensemble Approach
We propose using an ensemble of unsupervised similarity measures to improve the performance when assessing code similarity. The rationale behind combining several unsupervised similarity measures is to offer a more accurate evaluation than strategies using just one measure [27]. Therefore, we first define the problem, introduce the concept of an ensemble, and discuss bagging and boosting as the two most prominent techniques for building ensembles. These techniques for building ensembles help aggregate different unsupervised similarity measures, leading usually to an enhanced strategy.
@@ 3.1 Problem Statement
Some applications can be built from this definition [34]. For example, code clone detection can be performed by applying a threshold value to differentiate between clones and non-clones [23].
@@
In addition, something similar can also be done to determine whether or not the code was plagiarized in an academic environment [3], for example. These two challenges of a very similar nature will be studied in more detail later in the paper.
@@ 3.2 Ensemble Definition
There are several techniques for learning ensembles, almost all based on a purely data-driven strategy. In the context of this work, and for practical reasons, we focus on two: bagging and boosting. Other techniques, such as stacking, will not be considered here because they have already been addressed in similar problems in the past, such as [26]. Although an in-depth study of each of these techniques is beyond the scope of this study, we offer below a brief description of how these ensemble-building strategies operate.
@@ 3.3 Bagging
The idea behind bagging is to train multiple instances of the same model with different data subsets and aggregate their predictions [6]. This strategy takes a list of code fragments, the number of subsets to create, and unsupervised similarity measures as input. It then averages the results of applying all measures to each subset.
@@ 3.4 Boosting
The idea behind boosting is to train multiple models, with each model instance focuses on the errors of the previous one [4]. It is necessary to operate by sequentially applying a list of code fragments, the number of rounds to run, and a list of unsupervised measures as input. This sequential application allows for iterative refinement of the code similarity scores, with the weights being updated based on performance, leading to the final scores.
@@ 3.5 Differences with the Transformer architecture
Using an ensemble of unsupervised semantic similarity measures for code similarity assessment offers several advantages over relying solely on a model like CodeBERT [9], or its more powerful variant GraphCodeBERT [12]. The ensemble approach combines different unsupervised similarity measures, each capturing unique aspects of code, thus providing a diverse analysis that can reduce individual biases and improve performance.
@@ 4 Empirical Evaluation
In previous work [28], we show that unsupervised similarity measures exhibit strengths in different domains. Some excel in textual comparisons, making them ideal for identifying cloned text. Others prioritize the detection of functional similarity, while specific measures specialize in uncovering structural resemblances between code and textual elements. The optimal choice of measure ultimately depends on the composition of the benchmark dataset. However, it is not reasonable to ask a human expert to do that since it would require many hours of work, and the result would be prone to errors since the task is far from trivial. The good news is that bagging and boosting algorithms, with their data-driven strategies, will decide the inclusion and importance of each of the simple unsupervised similarity measures.
@@ 4.1 Baseline
In this study, we compare our ensembles using two strategies. First, a weak baseline consists of obtaining better results than each similarity measure considered individually. Secondly, a strong baseline that consists of the state-of-the-art conformed by CodeBERT [9], and its variant GraphCodeBERT [12].
@@ 4.2 Small-scale Dataset
Firstly, we use the IR-Plag dataset [^1] https://github.com/oscarkarnalim/sourcecodeplagiarismdataset [21]. This dataset was designed for benchmarking code similarity techniques. It contains code files deliberately crafted to simulate patterns of academic plagiarism. Although primarily intended for plagiarism detection, its characteristics align with our goals due to the overlap between plagiarized and cloned code, which involves replication (though the intent may differ). The IR-Plag dataset encompasses a variety of complexities without strictly classifying the clone types.
The dataset consists of seven original code files. A substantial portion, 355 files (77%), are labeled as plagiarized, indicating a high prevalence of duplication. 105 files are considered non-plagiarized, potentially representing modified or derivative works. This yields a total of 467 code files. Collectively, these files contain 59,201 tokens, with 540 unique tokens underlining the dataset’s lexical diversity. File sizes exhibit variation, ranging from 286 tokens (largest) to 40 tokens (smallest), with an average file size of approximately 126 tokens. Therefore, although small in size, it is a dataset with a high compositional diversity of programming elements.
@@ 4.3 Large-scale Dataset
Secondly, we use the BigCloneBench dataset, which is a benchmark collection created to evaluate how well different code clone detection strategies perform [^2] https://github.com/microsoft/CodeXGLUE/tree/main/Code-Code/Clone-detection-BigCloneBench. It is helpful to tackle the challenge of finding duplicate code. It includes a variety of programming languages, and its data comes from actual software and student projects. This mix makes it an interesting resource for evaluating clone detection in real-life situations. This dataset is usually used in training solutions to make code more maintainable.
@@ 4.4 Evaluation Criteria
For that reason, the community prefers other metrics to assess the precision and coverage of a clone detection strategy, such as:
  • Precision: This measures the accuracy of an approach by determining the fraction of true positive identifications among all detected code fragments. Higher precision indicates better reliability. It is defined as:
    Precision=True PositivesTrue Positives+False $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Positives}}$
  • Recall: It evaluates the completeness of an approach by measuring the fraction of true positive clones in the dataset that were correctly identified. Higher recall implies that fewer actual clones were missed.
    Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$
  • F-measure: This metric assesses the performance of an approach by harmonizing precision and recall, ensuring a balanced evaluation of both metrics.
    $F\text{-measure}=2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
This evaluation method is popular because it emphasizes the importance of positive classes by separately assessing false positives (precision) and false negatives (recall). It imposes penalties on the model for not identifying positive cases and for incorrect positive predictions. Subsequently, one can compute a harmonic mean of these two metrics, known as the F-measure, which serves as a basis for ranking different techniques.
@@ 4.5 Hardware requeriments
Please note that all the computations of similarity measures and the learning of their ensembles has been done on a CPU 11th Gen Intel(R) Core(TM) i7-1185G7 a 3.00GHz with 32GB for RAM (TDP of 50W approx.), while the configuration and execution of the baselines with CodeBERT and GraphCodeBERT has required a GPU Tesla V100-PCIE-16GB (TDP of 300W approx.).

### v2

# Softwarepark 32a, 4232 Hagenberg, Austria

@@ Abstract
Ensemble Learning, Similarity Measures, Code Similarity
@@ 1 Introduction
Automatically assessing the degree of similarity between source code fragments presents several challenges ranging from syntactic variability or the use of complex algorithmic structures to the employment of diverse abstraction levels across programming languages juergens2009code . While established techniques and tools exist to address this problem, significant limitations must be addressed. For example, the best-performing approaches need vast training data, their insights are practically intelligible to the human operator, and their associated carbon footprint during the training phase does not bring any positive societal outcome key-martinez-jfis .
In order to overcome these issues, this research proposes a novel ensemble approach for improved code similarity assessment. The method combines the outputs of multiple similarity measures to produce better similarity scores. This approach is based on the principle that the strengths of diverse similarity measures can complement each other and mitigate individual weaknesses, leading to improved systemic performance.
Extensive empirical evaluations on several datasets have demonstrated the effectiveness of our ensemble learning method. Our method outperformed traditional single-model approaches in all cases (both in scenarios with small and large datasets). Furthermore, our method achieved comparable performance to state-of-the-art methods when solving small datasets without reliance on deep learning techniques, making it a solution to consider under the circumstances such as the absence of significant amounts of training data, the need for interpretability, and the wish for a reduced carbon footprint when training.
Therefore, our contribution is not only to guide the aggregation of appropriate unsupervised similarity measures for code similarity but also to pave the way for applications such as source clone detection saini2018code or code plagiarism assessment novak2019source in realistic scenarios where there is not too much data for training. In this way, the major contributions of this study are:
  • We build upon previous work key-martinez-swqd to curate a comprehensive collection of unsupervised similarity measures to compare the likeness of different code fragments. This collection facilitates a systematic exploration of the most effective measures for assessing source code similarity at low computational cost.
@@
  • We evaluate the proposed ensembles focuses on metrics such as precision, recall, and f-measure. Our findings show that ensembles outperform individual similarity measures and have potential to rival even the most sophisticated state-of-the-art techniques in the absence of significant training data.
@@ 2 State-of-the-art
The existence of many programming languages and individual coding styles poses a challenge for determining the degree of likeness between source code fragments ain2019systematic . Despite that, unsupervised similarity measures demonstrate reasonable adaptability when working with code written in multiple languages or according to various styles, so it can be accepted that they are good indicators of code similarity key-martinez-swqd . This feature is also important in many scenarios, such as maintenance, whereby increased complexity is commonplace higo2002software .
The research community widely acknowledges that the emerging strategies based on language models offer superior performance in this context karmakar2021pre . However, our hypothesis is that unsupervised methods for code similarity evaluation offer several advantages for addressing key aspects that should not be overlooked:
  • Capturing the meaning of code is a complex task, as functionally equivalent code fragments may exhibit superficial differences or, on the contrary, share surface-level similarities while having entirely different purposes. Some unsupervised techniques specifically focus on the semantic intent rather than syntactic details.
  • Since unsupervised techniques eliminate the need for labeled training data, they reduce the overhead of establishing a ground truth. While labeled examples remain valuable for validating the effectiveness of these methods, these examples are not a prerequisite for deployment.
  • Codebases often contain comments and other non-functional elements embedded within the core logic. Specific unsupervised methods can distinguish between meaningful code patterns and just incidental material.
  • Unsupervised methods can accommodate the diversity of programming languages and coding styles. This mitigates the need for a complex strategy that attempts to account for all possible variations.
We acknowledge the superior capabilities of language models when confronted with this challenge. However, we also recognize and aim to demonstrate that alternative models, which excel in specific areas, could be employed under certain plausible conditions across a wide range of scenarios. For instance, these models may be more suitable in situations where only limited datasets are available or where the working logic of the proposed solution needs to be clearly understood.
@@ 2.1 Unsupervised Code Similarity
There are various techniques to assess the similarity between source code fragments, each with its unique perspective based on certain features of the code fragments being compared. In previous work key-martinez-swqd , we perform a categorization of exiting approaches consisting of twenty-one distinct unsupervised strategies. Nonetheless, according to the same research study key-martinez-swqd , it is unrealistic to think that these techniques alone can deliver performance equivalent to the state-of-the-art.
@@ 2.2 Supervised Code Similarity
Supervised code similarity is a well-studied field and has yielded the best results to date wang2020detecting ; wei2017supervised ; white2016deep ; zhang2019novel ; yu2019neural . Although among the existing approaches, CodeBERT key-codebert and its variant GraphCodeBERT guo2020graphcodebert stand out. CodeBERT is a variant of BERT key-Bert , and empirical results show that it is the leading technology to understand source code expressed in different general-purpose programming languages. It excels at tasks like generating documentation, creating code, and comparing similar pieces of code. The idea behind CodeBERT is to combine techniques from language processing with an understanding of code and training on a wide range of programming languages and their associated comments. Results show CodeBERT is very good at detecting similarities between source code fragments, as demonstrated by its performance in the resolution of benchmark datasets for clone detection key-codebert . However, an additional fine-tuning phase on the pre-trained model over the actual data is usually required to obtain optimal results.
@@ 2.3 Contribution over the state-of-the-art
This research introduces a novel ensemble learning approach for code similarity assessment. The rationale behind combining multiple unsupervised similarity measures is to benefit from the strengths of individual approaches while mitigating their weaknesses. This strategy is helpful in scenarios where deep learning is not optimal due to the absence of significant training data, when prioritizing model interpretability without significantly compromising performance, or even when seeking to minimize the environmental impact.
@@ 3 Ensemble Approach
We propose using an ensemble of unsupervised similarity measures to improve the performance when assessing code similarity. The rationale behind combining several unsupervised similarity measures is to offer a more accurate evaluation than strategies using just one measure key-martinez-ijseke . Therefore, we first define the problem, introduce the concept of an ensemble, and discuss bagging and boosting as the two most prominent techniques for building ensembles. These techniques for building ensembles help aggregate different unsupervised similarity measures, leading usually to an enhanced strategy.
@@ 3.1 Problem Statement
Some applications can be built from this definition roy2009comparison . For example, code clone detection can be performed by applying a threshold value to differentiate between clones and non-clones karnalim2021explanation .
@@
In addition, something similar can also be done to determine whether or not the code was plagiarized in an academic environment aniceto2021source , for example. These two challenges of a very similar nature will be studied in more detail later in the paper.
@@ 3.2 Ensemble Definition
There are several techniques for learning ensembles, almost all based on a purely data-driven strategy. In the context of this work, and for practical reasons, we focus on two: bagging and boosting. Other techniques, such as stacking, will not be considered here because they have already been addressed in similar problems in the past, such as key-martinez-mlwa . Although an in-depth study of each of these techniques is beyond the scope of this study, we offer below a brief description of how these ensemble-building strategies operate.
@@ 3.3 Bagging
The idea behind bagging is to train multiple instances of the same model with different data subsets and aggregate their predictions breiman1996bagging .
Formally, let $D$ denote the original dataset and $M$ the number of subsets to create. The subsets $D_{i}$ for $i\in\{1,2,\ldots,M\}$ are generated through random sampling with replacement. For each subset $D_{i}$, a model $f_{i}$ is trained. Given an unsupervised similarity measure $S$, we compute the similarity score for each subset:
$S_{i}=S(f_{i},D_{i})\quad\text{for }i\in\{1,2,\ldots,M\}.$
The final aggregated prediction is obtained by averaging the individual predictions from all models:
$\hat{y}=\frac{1}{M}\sum_{i=1}^{M}f_{i}(D).$
This technique helps reduce variance and improve the stability of the model.
@@ 3.4 Boosting
The idea behind boosting is to train multiple models, with each model instance focusing on the errors of the previous one bentejac2021comparative .
Formally, let $D$ denote the dataset and $N$ the number of boosting rounds. We start with an initial model $f_{1}$ and iterate sequentially through each round. For each round $t\in\{1,2,\ldots,N\}$, we train a model $f_{t}$ with a focus on the errors made by the previous models. Let $S$ denote a set of unsupervised similarity measures.
At each round, we update a weight distribution $w_{t}$ over the dataset based on the errors of $f_{t}$, aiming to emphasize the misclassified cases. The similarity scores for the current model $f_{t}$ are calculated as:
$S_{t}=S(f_{t},D)\quad\text{for }t\in\{1,2,\ldots,N\}.$
The final aggregated prediction is obtained by combining the models using their respective weights:
$\hat{y}=\sum_{t=1}^{N}\alpha_{t}f_{t}(D),$
where $\alpha_{t}$ is the weight assigned to each model $f_{t}$ based on its performance in reducing the error. This sequential training allows for iterative refinement of the predictions.
@@ 3.5 Differences with the Transformer architecture
Using an ensemble of unsupervised semantic similarity measures for code similarity assessment offers several advantages over relying solely on a model like CodeBERT key-Bert , or its more powerful variant GraphCodeBERT guo2020graphcodebert . The ensemble approach combines different unsupervised similarity measures, each capturing unique aspects of code, thus providing a diverse analysis that can reduce individual biases and improve performance.
@@ 4 Empirical Evaluation
In previous work key-martinez-swqd , we show that unsupervised similarity measures exhibit strengths in different domains. Some excel in textual comparisons, making them ideal for identifying cloned text. Others prioritize the detection of functional similarity, while specific measures specialize in uncovering structural resemblances between code and textual elements. The optimal choice of measure ultimately depends on the composition of the benchmark dataset. However, it is not reasonable to ask a human expert to do that since it would require many hours of work, and the result would be prone to errors since the task is far from trivial. The good news is that bagging and boosting algorithms, with their data-driven strategies, will decide the inclusion and importance of each of the simple unsupervised similarity measures.
@@ 4.1 Baseline
In this study, we compare our ensembles using two strategies. First, a weak baseline consists of obtaining better results than each similarity measure considered individually. Secondly, a strong baseline consists of the state-of-the-art conformed by CodeBERT key-Bert and its variant GraphCodeBERT guo2020graphcodebert .
@@ 4.2 Small-scale Dataset
Firstly, we use the IR-Plag dataset key-karnalim . This dataset was designed for benchmarking code similarity techniques. It contains code files deliberately crafted to simulate patterns of academic plagiarism. Although primarily intended for plagiarism detection, its characteristics align with our goals due to the overlap between plagiarized and cloned code, which involves replication (though the intent may differ). The IR-Plag dataset encompasses a variety of complexities without strictly classifying the clone types.
The dataset consists of seven original code files. A substantial portion, 355 files (77%), are labeled as plagiarized, indicating a high prevalence of duplication. One hundred five files are considered non-plagiarized, potentially representing modified or derivative works. This yields a total of 467 code files. Collectively, these files contain 59,201 tokens, with 540 unique tokens underlining the dataset’s lexical diversity. File sizes exhibit variation, ranging from 286 tokens (largest) to 40 tokens (smallest), with an average file size of approximately 126 tokens. Therefore, although small in size, it is a dataset with a high compositional diversity of programming elements.
@@ 4.3 Large-scale Dataset
Secondly, we use the BigCloneBench dataset, which is a benchmark collection created to evaluate how well different code clone detection strategies perform. It is helpful to tackle the challenge of finding duplicate code. It includes a variety of programming languages, and its data comes from actual software and student projects. This mix makes it an interesting resource for evaluating clone detection in real-life situations. This dataset is usually used in training solutions to make code more maintainable.
@@ 4.4 Evaluation Criteria
For that reason, the community prefers other metrics to assess the fitness a clone detection strategy. For instance, the Precision/Recall evaluation method is popular because it emphasizes the importance of positive classes by separately assessing false positives (precision) and false negatives (recall). It penalizes the model for not identifying positive cases and for incorrect positive predictions. Subsequently, one can compute a harmonic mean of these two metrics, known as the F-measure, which serves as a basis for an homogeneous ranking.
@@ 4.5 Hardware requeriments
Please note that all the computations of similarity measures and the learning of their ensembles have been done on a CPU 11th Gen Intel(R) Core(TM) i7-1185G7 a 3.00GHz with 32GB for RAM (TDP of 50W approx.), while the configuration and execution of the baselines with CodeBERT and GraphCodeBERT has required a GPU Tesla V100-PCIE-16GB (TDP of 300W approx.).